OpenStudy (anonymous):

solve for x. 4x(x-5)-5x(x-4)=-1

OpenStudy (anonymous):

First distribute. Then combine like terms. Then solve the quadratic.

OpenStudy (anonymous):

ok, 4x-20-5x+20= -1?

OpenStudy (anonymous):

Nope. \[4x(x-5) - 5x(x-4) = -1\]\[4x(x) - 4x(5) - 5x(x) -5x(-4) = -1\]

OpenStudy (anonymous):

and continue from there. You have to distribute the 4x to each term.

OpenStudy (anonymous):

unless you mean for the x to be multiplication symbol and not a variable.

OpenStudy (anonymous):

no the x is not multiplication symbol.... ok im going to try it, ty for the advice i really appreciate it.

OpenStudy (anonymous):

Just remember how to distribute: \[a(b+c) = a(b) + a(c)\] That's all I did above, just with different values.

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

ok, x = [20 + -√400-4x25)]/2 ????

OpenStudy (anonymous):

x =[20 + -√300)] /2

OpenStudy (anonymous):

im trying to work it, ill get it im sure.

OpenStudy (anonymous):

Not quite right. What did you get for the quadratic after you combined like terms??

OpenStudy (anonymous):

x^2+20x - 25 = 0

OpenStudy (anonymous):

thanks for the help. ill continue to work at it. it has been over 25 yrs since schooling.